# Can Graph Learning Improve Task Planning?

### v1

@@ 7 Conclusions
This paper presents an initial exploration into graph-learning-based approaches for task planning. Through theoretical analysis, we demonstrate the inductive bias of the attention mechanism and the utility of auto-regressive loss impedes their effectiveness in task planning. We propose to integrate GNNs for task graph analysis, which yields performance improvements across a range of LLMs and datasets.
@@
Limitations: Despite the encouraging performance, there are limitations that highlight significant opportunities for enhancement. Firstly, our proposed method, while effective, is straightforward; more sophisticated GNN-based decision-making algorithms could potentially offer further improvements. Secondly, in our current framework, GNNs function as an external module with limited interaction with LLMs. Exploring the synergies between GNNs and LLMs, particularly by incorporating GNN outputs as tokens within LLMs, presents an interesting avenue for research. Thirdly, the construction of the task graph currently requires manual effort. Investigating automated graph generation techniques for this application is another promising direction for future work.
Appendix A More Related Works
A.1 LLMs for Graphs
@@
With the breakthroughs in LLMs, there has been a surge of interest in applying LLMs to graph-related problems. GPT4Graph [10] and NLGraph [41] are two prominent benchmarks designed to evaluate the performance of LLMs in the context of graph tasks. They encompass a wide spectrum of challenges, various input formats, and state-of-the-art prompting techniques, demonstrating that LLMs possess basic graph processing capabilities. Importantly, the choice of prompts and formats significantly influences performance. However, these benchmarks also expose the models’ susceptibility to spurious correlations within graphs. For instance, GPT-4 achieves only about $50\%$ accuracy on shortest-path tasks, even with the use of complex prompts. GraphInstruct [24] attempts to fine-tune LLMs on graph-theory-related tasks, resulting in improved performance, though it remains far from satisfactory. Despite these empirical efforts, there is a limited theoretical understanding of these evaluation results. The analysis in Section 3.2 aims to shed light on the empirical observations reported in these studies.
@@
Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [37] and GraphLLM [6]. These approaches have demonstrated significant improvements in performance on GNN-related tasks. However, they have not yet been applied to task planning due to the lack of extensive training data. A promising future direction involves using task planning data generated by GPT to fine-tune graph foundation models, such as GraphGPT [37], and applying them to task planning.
A.2 Theoretical Analysis of Reasoning
@@
A.3 GNNs for Decision-making
@@
Appendix B Proofs
B.1 Proof of Theorem 1
Assumption 1.
Each function $f,\oplus$ in (2) can be approximated by constant size MLP.
Assumption 2.
The aggregation function $\square$ in (2) is one of min, max, sum, mean.
Theorem 3.
(Expressiveness) Assume the input format is given in (3) and $f,g,\square$ in DP update (2) satisfy the assumptions 1 and 2. There exists a log-precision constant-depth and constant-width Transformer that simulates $k$ steps of DP update in (2). As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
Proof.
Block 1 - Initial State Broadcast: The goal of the first block is to broadcast the initial states from the initial state token to node tokens. (1) Use MLPs to recover the digits of the answer tokens and put them in the first placeholder if $e^{type}_{k}==0$; (2) Copy the first placeholder from answer token to its previous node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos1}}_{k}-e^{\text{pos1}}_{j})^{2}<\delta\}$; (3) Broadcast the first placeholder with MEAN in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{type1}}_{k}-e^{\text{type1}}_{j})^{2}+(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}]]$.
Block 2 - Edge Feature Operations: The goal of the second block is to copy the edge features from the edge feature token to the corresponding node tokens. (1) Use MLPs to recover the digits of the edge feature tokens and put them in the second placeholder if $e^{\text{type1}}==2$; (2) Copy the second placeholder from the edge feature token to the node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos2}}_{k}-e^{\text{pos2}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}]]$.
Block 3 - Message Preparation: (1) Use MLPs to compute $\oplus$ and place the results in the third placeholder. Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}],\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{i}]]$; (2) Use MLPs to clean up the first and second placeholder. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos1}},\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{i}]]$
Block 4 - Message Passing: The goal of the fourth block is to compute $\square$ and $f$. (1) Use one or two attention heads (one for max, min, mean aggregations, and two for sum aggregations) to perform the aggregation operation. This is achieved by using MEAN or MAX or SUM for the first placeholder and setting $\mathcal{S}_{k}=\{j|(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}+(e^{\text{type2}}_{k}-e^{\text{type2}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\square_{v_{j}\in\mathcal{T}(u_{i})}\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{j}]]$; (2) Use MLPs to compute $f$. Now the state for every node token is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},f(\square_{v_{j}\in\mathcal{T}(i)}(\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{j}]))]$.
After four blocks, the final state for every node token $u_{i}$ is given by $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[1][u_{i}]]$. $\text{Answer}[k][u_{i}]$ can be obtained by repeating the above four blocks $k$ times. ∎
Lemma 1.
B.2 Proof of Proposition 2
Proposition 2.
Proof.
B.3 Proof of Theorem 2
Theorem 4.
Proof.
$-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\bm{v}_{i,j,u}\log\hat{\bm{v}}_{i,j,u}=-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\mathbb{I}_{u=v_{i,j}}\log\hat{\bm{v}}_{i,j,u}\overset{(a)}{=}-\sum_{t,v_{j-1},u}N_{t,v_{j-1},u}\log\hat{\bm{v}}_{i,j,u}$
where (a) uses the assumption that the output logits are determined by target node $t$ and the current node $v_{i-1}$. In (b), we assume that $\sum_{u}N_{t,v_{i-1},u}\neq 0$. If $\sum_{u}N_{t,v_{i-1},u}=0$, then the corresponding logits will not affect the loss function and $\hat{\bm{v}}_{i,j,u}$ can take any number. The cross-entropy is minimized when $\hat{\bm{v}}_{i,j,u}=\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}$. ∎
Appendix C More Discussions
C.1 Examples of Dynamic Programming
where $\mathcal{T}(i)=\{i\}\cup\{j\;|j<i\;\text{and}\;\text{arr}[j]<\text{arr}[i]\}$ denotes the set of states that can transfer to state $i$, the aggregation function $\square$ is implemented as $\max(\cdot)$, and the cost $c[i][j]$ is $1$ for those candidate states that are not equal to state $i$ as adding the element leads to a longer subsequence while 00 for the state itself.
Bellman-Ford Algorithm: The Bellman-Ford algorithm is also a classic dynamic programming algorithm used to find the shortest path from a single source vertex to all other vertices in a weighted graph. The core idea behind the Bellman-Ford algorithm is that the distance from the source vertex to a target vertex can be computed as the minimum distance from the source to any of the target’s neighboring vertices, plus the weight of the edge connecting the neighbor to the target. Therefore, the state transition function for the Bellman-Ford algorithm can be expressed as:
where $\text{Answer}[k][i]$ represents the length of the shortest path from the source vertex to node $i$ at the $k$-th iteration, $\mathcal{T}(i)=\mathcal{N}^{-}(i)$ denotes the set of in-neighbors of node $i$, and $w[j][i]$ denotes the weight from node $j$ to $i$. The aggregation function $\square$ is implemented as $\min(\cdot)$ as we try to find the shortest path.
Travelling Salesman Problem: This problem is defined as, given a set of cities and the distances between every pair of cities, finding the shortest possible route that visits every city exactly once and returns to the starting point. If we regard the set of already visited city $\mathcal{S}$ ending at $i$-th city as the current state, then states that can transfer to current state are those that ending city can reach $k$. Therefore, the state transition function for the TSP problem can be expressed as:
where $\text{Answer}[k][i][\mathcal{S}]$ represents the cost of the shortest tour that visits all the cities in the set $\mathcal{S}$ and ends at the $i$-th city, the aggregation function is still implemented as $\min(\cdot)$ since we aim to find the shortest path, and $w[j][i]$ denotes the distance from city $j$ to city $i$.
C.2 Permutation Invariance Test of LLMs
Figure 4: Illustrative Examples of LLMs Failure to Solve Graph Problems under Node-ID Permutation. Experiments were repeated for 30 times.
Appendix D Implementation Details
D.1 Details of Prompt Templates
To facilitate reproducibility, we provide prompt templates for all LLM-related queries, including direct inference and GraphSearch in Tables 3 and 4, respectively. GraphSearch consists of two types of LLM queries: Task Assessment, where LLM evaluates the suitability of each candidate task to address current step’s requirement during the iterative search, and Path Selection, where LLM selects the optimal path to fulfill user’s request after the search.
@@ D.1 Details of Prompt Templates
D.2 Implementation of Training-free Methods
In this subsection, we present the implementation details of two training-free baselines, including LLM’s direct inference and GraphSearch. Method illustrations are shown in Figure 5.
Figure 6: Illustration of our GNN-enhanced Task Planning. We first prompt LLM to decompose complex and implicit user request into several manageable sub-tasks. We then use either a training-free or training-required GNN to sequentially align each step description with a suitable task, ultimately generating the invocation path.
LLM’s Direct Inference: During experiments, we uniformly apply 1-shot in context learning for LLM’s direct inference of task invocation path. For open-sourced LLMs, the temperature parameter is set to 0.2.
GraphSearch: This algorithm conducts an iterative search on the task graph to identify an optimal task invocation path that can best satisfy a given request. In each iteration, the neighbors of the last selected task are considered as candidates. These candidates are evaluated by LLM for their suitability for the current step. The search process follows a depth-first approach. After the task assessment in the final step, a set of potential invocation paths is generated. Subsequently, LLM is prompted to select the most appropriate path from these options. The GraphSearch algorithm is implemented in three distinct variants, each employing a unique task selection strategy:
SGC: For the implementation of SGC, the parameter $\alpha$ is uniformly set to 0.7. Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC on HuggingFace dataset, Roberta-large-355M [26] serves as the text encoder. For all other datasets and LLMs, the e5-355M [44] configuration is employed.
D.3 Implementation of Training-based GNNs
LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-355M [44] as the LM backbone. For the graph encoder, our setup includes a single layer with a hidden dimension of 1024. During the model training, we set the batch size to 512 and run for 20 epochs with a learning rate of 1e-3. We use the Adam optimizer [17] and implement an early stopping mechanism with a patience of 5 epochs to prevent over-fitting. All experiments are conducted on a single NVIDIA A100-80G GPU.
Training Data Preparation: From each dataset in TaskBench, we randomly select $3,000$ samples to create the trainset. The original data includes specific task steps and corresponding task invocation paths. Therefore, we first employ a topological sort to align each task step accurately with corresponding task, forming “<step, ground-truth task>” pairs. Then, for each pair, we randomly sample 2 negative tasks to constitute the “<step, positive task, negative task>” triplets for model training. These negative samples are selected based on how textually similar they are to the positive one, creating a robust differentiation challenge for the model.
Choices of Different Configurations: Our training-based method offers two configuration options: training only the GNN while keeping the LM frozen, or co-training both the LM and GNN. The first configuration is designed to explore GNN’s capability in task retrieval. The latter leverages LM’s dataset-specific semantic embeddings to enhance performance. Illustrations of these configurations are provided in Figure 7. For the co-training setup, we use a learning rate of 2e-5 and a training duration of 10 epochs.
D.4 Fine-tuning LLMs
To explore the effectiveness of our proposed framework on fine-tuned LLMs, we employ Supervised Fine-Tuning (SFT) on the CodeLlama-7B and Vicuna-13B models. We use the same set of $3,000$ samples from GNN training as the LLMs’ fine-tuning data. In fine-tuning the LLM with LoRA [13], we set the lora_r parameter (dimension for LoRA update matrices) to 8 and the lora_alpha (scaling factor) to 16. The dropout ratio is set to 0.1, the batch size to 2, and we conduct training over 2 epochs with a learning rate of 1e-5. For HuggingFace dataset, the maximum input length is set to 800, while the maximum output length is 400. For Multimedia and Daily Life datasets, which contain a larger number of tasks and require longer textual inputs, we set the maximum input and output length to 1000 and 500, respectively. We utilize 2 NVIDIA A100-80G GPUs for fine-tuning the LLMs. Each dataset undergoes specific fine-tuning for both the CodeLlama-7B and Vicuna-13B models.
Appendix E Datasets
E.1 Overview
@@
We provide the statistics of experimental datasets from two task planning benchmarks in Table 5. The illustrative examples from each dataset are shown in Figure 8. For datasets from TaskBench, each sample consists of original user request, corresponding decomposed task steps, and ground-truth task invocation path. As RestBench includes only user requests and corresponding API invocation sequences, we prompt GPT-4 to infer decomposed task steps aligned with each invoked API, thereby finalizing the dataset.
E.2 Reformatting Details of RestBench
@@
The TMDB dataset from RestBench, focuses on movie-related searching and recommending functions. To align RestBench with our experiments, we have implemented the following processing steps: Reformatting original APIs by assigning unique task names and descriptions: APIs in RestBench were represented by request paths, such as “GET /movie/top_rated”, referring to the API that retrieves top-rated movies on TMDB. To enhance semantic differentiation among APIs, we first prompt GPT-4 to assign a unique name and a detailed functional description to each API. These names and descriptions were then manually verified and refined. For example, the API previously mentioned is renamed “Get Top-Rated Movies” with the description “This API retrieves a list of the highest-rated movies.” Note that though the original TMDB dataset contains 54 APIs, some were never invoked in any data examples. Therefore, we focus only on those APIs that appeared in at least one user request, resulting in a refined set of 46 APIs.

### v2

@@ 6 Conclusions
This paper presents an initial exploration into graph-learning-based approaches for task planning in language agents. Through theoretical analysis, we demonstrate the inductive bias of the attention mechanism and the utility of auto-regressive loss impedes their effectiveness in task planning. We propose to integrate GNNs for task graph analysis, which yields performance improvements across a range of LLMs and planning benchmarks.
@@
Limitations: Despite the encouraging performance, there are limitations that highlight significant opportunities for enhancement. Firstly, our proposed method, while effective, is straightforward; more sophisticated graph-learning-based decision-making algorithms could potentially offer further improvements. Secondly, the construction of the task graph currently requires manual effort. Investigating automated graph generation techniques for diverse applications is another promising direction for future work.
Acknowledgements
This work is partly supported by grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14217622).
Appendix A Related Works and Discussions
A.1 Planning Algorithms in LLMs
The existing studies of task planning approaches can be categorized into several directions, including task decomposition, multi-plan selection, the use of external planners, reflection, and memory-aided planning [21]. Task decomposition methods, such as the chain-of-thought approach [62], employ the divide-and-conquer strategy, utilizing LLMs for both task decomposition and sub-task planning. The application of this method to task planning is detailed in Section 2.2 and is referred to as “Direct” in the baseline comparison. Multi-plan selection strategies, exemplified by the tree-of-thought [70] and graph-of-thought [4], leverage search-based methods to generate plans. Subsequently, LLMs evaluate these plans to select the most effective one. The “GraphSearch” methods used in our baselines fall into this category. External planner approaches [30] use LLMs to convert the problem into Planning Domain Definition Language (PDDL) and then employ classic solvers to address the planning problem. PDDL requires a pre-defined goal, for example, moving the blocks from one state to another state. However, the goal of task planning investigated in language agents deals with more flexible and personal goals, spanning personal needs in video, text, and image processing. Translating these goals into formal PDDL is very difficult. We instead demonstrate that GNNs can serve as an effective external planner in this application. Reflection-based methods [47] focus on reflecting upon experiences to refine the plan, while memory-aided planning approaches [72] utilize external experiences, such as those from search engines. These approaches are deployed in interactive environments and orthogonal to this paper.
A.2 Task Planning in Traditional AI
Apart from task planning in language agents, there is also a domain in traditional AI called task planning [1, 6, 17, 54, 43, 50, 34, 10]. Task planning in traditional AI is defined as $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{C},\mathcal{G},s_{0})$, where $\mathcal{S}$ is the states, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, a cost function $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$, a set of goal states $\mathcal{G}\subseteq\mathcal{S}$, and an initial state $s_{0}$. An agent following a policy $\pi:\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ will start in state $s_{0}$, then repeatedly choose an action $a\sim\pi(a|s)$ and execute it to reach a new state $s^{\prime}\sim\mathcal{T}(s^{\prime}|s,a)$, incurring a cost $\mathcal{C}(s,a)$ along its way. An optimal policy $\pi^{*}(a|s)$ is one that reaches the goal state with probability $1$ while minimizing the total expected cost. Traditionally, task planning is solved by reinforcement learning approaches [6] and heuristic A approaches [17]. The neural network-based approaches are employed to accelerate the computation and improve the performance [54, 43, 50, 34, 10].
The task planning in language agents is a different application. It cannot be solved as a constraint satisfaction problem, since both the features of task graph and user request are expressed in natural language. Compared with traditional planning, task planning for language agents involves diverse and open-ended goals due to the varied personal requirements users have. For example, on platforms like Hugging Face, users’ intentions span across video, text, and image domains. On the contrary, classic planning has a fixed goal for a given domain, e.g., in the n-puzzle [54], the goal is formally as placing the tiles in numerical order. Within this new application domain, while existing research primarily focuses on prompt design for pre-trained LLMs, our work underscores the importance of traditional planning methods, such as GNNs, in complementing LLMs.
A.3 Planning in Agents and Neuroscience
Planning is a pivotal topic in both agents and neuroscience, where graphs play an indispensable role. We believe the concepts and insights presented in this paper are useful to these fields.
The TaskBench, RestBench, and UltraTool dataset used in this paper belongs to the tool agents. The Math agent, AlphaGeometry, employs LLMs to generate auxiliary constructions in geometry [55]. Considering lemmas as nodes and their interdependencies as edges, the endeavor to prove a theorem resembles the task of identifying a route to the theorem node within the graph constituted by potential lemma nodes and the edges that represent their interdependencies. There are no explicit task graphs in game agents [57], embodied agents [20], and code agents [47]. The core strategy in these domains is to employ verbal reinforcement learning within LLMs. The state and transitions in reinforcement learning can be modeled as nodes and edges in the graph. In addition, there are case-by-case graph models in these agent applications. For example, in code agents, one can view the class as the nodes and dependencies as the edges. In the embodied agents, the objects in the environment can be viewed as the nodes.
In the neuroscience, animal planning is often assessed through path planning in mazes [2, 63]. Inspired by these animal experiments, planning testbenches have been developed for LLMs [36]. A computational model known as the Tolman-Eichenbaum Machine (TEM) has been proposed to decipher the mechanisms of general planning in animals across various environments, such as mazes [63]. The TEM model posits that hippocampal cells, including place and landmark cells, remap between environments, while entorhinal cells exhibit a range of properties that mirror spatial responses, including grid, band, border, and object-vector cells. In essence, hippocampal cells map sensory inputs onto locations in abstract graphs and remap, and entorhinal cells execute graph operations.
A.4 LLMs for Graphs
@@
With the breakthroughs in LLMs, there has been a surge of interest in applying LLMs to graph-related problems [26, 27]. GPT4Graph [14] and NLGraph [58] are two prominent benchmarks designed to evaluate the performance of LLMs in the context of graph tasks. They encompass a wide spectrum of challenges, various input formats, and state-of-the-art prompting techniques, demonstrating that LLMs possess basic graph processing capabilities. Importantly, the choice of prompts and formats significantly influences performance. However, these benchmarks also expose the models’ susceptibility to spurious correlations within graphs. For instance, GPT-4 achieves only about $50\%$ accuracy on shortest-path tasks, even with the use of complex prompts. GraphInstruct [33] attempts to fine-tune LLMs on graph-theory-related tasks, resulting in improved performance, though it remains far from satisfactory. Despite these empirical efforts, there is a limited theoretical understanding of these evaluation results. The analysis in Section 3.3 aims to shed light on the empirical observations reported in these studies.
@@
Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [52], GraphLLM [9], and GraphToken [37]. These approaches have demonstrated significant improvements in performance on GNN-related tasks. However, they have not yet been applied to task planning due to the lack of extensive training data. A promising future direction involves using task planning data generated by GPT to fine-tune graph foundation models, such as GraphGPT [52], and applying them to task planning. This paper proposes to use task planning as a new benchmark for this line of research.
A.5 Theoretical Analysis of Reasoning
@@
A.6 GNNs and GraphSearch for Combinatorial Optimization
@@
Appendix B Prompts
@@
Table 6: Prompt template for LLM’s filling in invocation parameters [44]
# GOAL #
Given a # USER REQUEST # and # PLANNED TASKS # to be invoked in sequence to solve this request, please fill up each invoked task’s invocation parameters.
The format must be in strict JSON format, like:
[…]
# REQUIREMENTS #
1. Consider each task’s input and output requirements, and carefully fill in the arguments for each task.
2. Analyze the resource dependencies, keeping in mind that these tasks are invoked sequentially to address the original request.
3. The number of predicted task_nodes must strictly align with the provided tasks.
# USER REQUEST #
{{ user request }}
# PLANNED TASKS #
{{ a list of previously GNN retrieved tasks }}
# DETAILS OF TASKS #
{{ details, i.e., input and output requirements of each planned task }}
# EXAMPLE #
{{ in-context learning examples }}
Now, please generate your response in a strict JSON format: # RESULT #
Appendix C Datasets
C.1 Overview
@@
We provide the statistics of experimental datasets from three task planning benchmarks in Table 7. The illustrative examples from each dataset are shown in Figure 4. For datasets from TaskBench [44], each sample consists of original user request, corresponding decomposed task steps, and ground-truth task invocation path. As RestBench [49] and UltraTool [19] include only user requests and corresponding API invocation sequences, we prompt GPT-4 to infer decomposed task steps aligned with each invoked API, thereby finalizing the dataset.
C.2 Reformatting Details of RestBench
The TMDB dataset from RestBench, focuses on movie-related searching and recommending functions. To align RestBench with our experiments, we have implemented the following processing steps:
@@ C.2 Reformatting Details of RestBench
Reformatting original APIs by assigning unique task names and descriptions: APIs in RestBench were represented by request paths, such as “GET /movie/top_rated”, referring to the API that retrieves top-rated movies on TMDB. To enhance semantic differentiation among APIs, we first prompt GPT-4 to assign a unique name and a detailed functional description to each API. These names and descriptions were then manually verified and refined. For example, the API previously mentioned is renamed “Get Top-Rated Movies” with the description: “This API retrieves a list of the highest-rated movies.” Note that though the original TMDB dataset contains $54$ APIs, some were never invoked in any data examples. Therefore, we focus only on those APIs that appeared in at least one user request, resulting in a refined set of $46$ APIs.